OpenStudy (anonymous):

1 to answer get a medal 14+6a-8=18

jimthompson5910 (jim_thompson5910):

14+6a-8=18 6a+6=18 6a=18-6 6a=12 a=12/6 a=2

OpenStudy (anonymous):

a=2

OpenStudy (anonymous):

i mean it = -2

OpenStudy (anonymous):

(26-14)/6=a

OpenStudy (anonymous):

a=2